# When is the Carathéodory metric a Kähler metric?

### v1

@@ Abstract
We prove that if the Carathéodory metric on a strictly pseudoconvex domain with a smooth boundary is a Kähler metric, then the domain is biholomorphic to a ball. Additionally, we highlight the relationships between the Lu constant and the holomorphic Bergman curvature.
footnotetext: Keywords. Carathéodory metric, Bergman metric, Kähler metric, Kähler-Einstein metric, biholomorphic map, holomorphic sectional curvature, Lu constant, strictly pseudoconvex domain Mathematics Subject Classification. Primary 32F45; Secondary 32T15, 32H02
@@ 1 Introduction
Question 1.1.
What are the domains where the Bergman, Carathéodory, Kähler-Einstein, and Kobayashi metrics are not pairwise distinct (up to scaling)?
Question 1.1 relates to a general problem raised by Yau [45, pp. 679] (in a slightly different form) on the characterization of manifolds whose Bergman metrics are Kähler-Einstein. Cheng and Yau [9] showed the existence of a complete Kähler-Einstein metric on a smoothly bounded strictly pseudoconvex domain in $\mathbb{C}^{n}$. Mok and Yau [33] later demonstrated that there is a complete Kähler-Einstein metric for any bounded pseudoconvex domain without boundary regularity conditions. For smoothly bounded strictly pseudoconvex domains, Cheng conjectured that the Bergman metric is Kähler-Einstein if and only if the domain is biholomorphic to a ball. Following the earlier works of Fu and the third author [19] and Nemirovski and Shafikov [36], this conjecture was confirmed by Huang and Xiao in [25]. Additionally, see the recent work [38] by Savale and Xiao for finite type domains in $\mathbb{C}^{2}$.
A famous theorem of Lu [31] says that a bounded domain in $\mathbb{C}^{n}$ with a complete Bergman metric of constant holomorphic sectional curvature must be biholomorphic to $\mathbb{B}^{n}$. In [16, 17], the first and third authors extended Lu’s theorem to the Bergman-incomplete situation and provided a multidimensional generalization of Carathéodory’s theorem on the continuous extension of biholomorphisms up to the closures. See also the recent work [24] by Huang and Li.
@@ 2. $2$
    where $t$ is measured with respect to the Poincaré metric on $\mathbb{D}$;
@@
In [10, 11], Cheung and the third author proved that if a bounded convex domain in $\mathbb{C}^{n}$ has a Hermitian metric of constant negative holomorphic sectional curvature, such that all metric components blow up to infinity at the boundary, then the domain is biholomorphic to $\mathbb{B}^{n}$.
In this note, we investigate the bounded domains in $\mathbb{C}^{n}$ where the Carathéodory metric coincides up to scaling with the Bergman or Kähler-Einstein metric, or more generally, is a Kähler metric. Our work is inspired by a recent paper by Gaussier and Zimmer [20], who used the Kobayashi metric to characterize domains whose universal coverings are biholomorphic to a ball (see the theorem below); they also obtained results concerning locally symmetric metrics.
@@ 4. $(d)$
For strictly pseudoconvex domains, we obtain the following main result, which is a Carathéodory-analogue of both Lu’s uniformization theorem and the resolution of the Cheng conjecture.
@@ 1. $(1)$
    The Carathéodory metric on $\Omega$ is a Kähler metric;
@@ 1. $(i)$
    By the remarkable works of Lempert [28, 27, 29], the Carathéodory metric and the Kobayashi metric coincide on convex domains. Theorem 1.2/1.3 implies that the Carathéodory metric on a bounded convex domain is a Hermitian metric if and only if the domain is biholomorphic to the unit ball. In particular, let $\mathcal{P}$ be a small generic perturbation of $\mathbb{B}^{n}$ such that $\mathcal{P}$ is convex in $\mathbb{C}^{n}$ but not biholomorphic to $\mathbb{B}^{n}$ when $n\geq 2$ (see the papers of Burns, Shnider and Wells [7], and Greene and Krantz[22]). Then, the Carathéodory metric on $\mathcal{P}$ is not a Hermitian metric when $n\geq 2$.
@@ 3. $(iii)$
The following significant fact, that the Bergman metric always dominates the Carathéodory metric on bounded Euclidean domains, is an immediate consequence of an inequality due to Lu.
Theorem 1.6 ([30], cf. [8]).
$B_{\Omega}(z,v)\geq C_{\Omega}(z,v),$
The above inequality is not sharp. Therefore, we consider the following Lu constant.
[…]
The definition of $L(\Omega)$ implies that
$L(\Omega)B_{\Omega}\geq C_{\Omega}.$ (1.1)
Theorem 1.7.
${B}\geq 2c^{2}_{B};$
$B(z_{0})=2c^{2}_{B}(z_{0}).$
Since a bounded pseudoconvex domain with a spherical boundary in $\mathbb{C}^{n}$, $n\geq 2$, is universally covered by a ball by the uniformization theorem (see [36, Theorem A.2]), we obtain the following corollary relating to the examples of Burns and Shnider [6] using the aforementioned methods:
Corollary 1.8.
Lastly, we extend a previous result [11] by Cheung and the third author from bounded convex domains to strictly pseudoconvex domains.
@@
Besides [20], our proofs of Theorems 1.5 and 1.9 depends heavily on the work of Bracci, Fornæss, and Wold [3]. (For some related observations, see an earlier work by Huang [23].) One of the key ideas is that at a point sufficiently close to the boundary of a smoothly bounded strictly pseudoconvex domain, there is a nonempty, open set of directions where the Kobayashi and Carathéodory metrics coincide. We observe that the Carathéodory metric has constant negative holomorphic sectional curvature in these directions, and in fact, has constant negative holomorphic sectional curvature near the boundary, if the metric is Kähler.
The organization of the paper is as follows. In Section 2, we give a proof of our main result, Theorem 1.5. In Section 3, we provide proofs of Theorems 1.7 and 1.9.
2 Proof of the Main Result
@@
Theorem 2.2 ([3, 20, 23]).
Suppose that $\Omega\subset\mathbb{C}^{n},n\geq 2,$ is a bounded strictly pseudoconvex domain with $C^{2}$ boundary. Then there exists $\epsilon>0$ such that
${C_{\Omega}(z,v)}={K_{\Omega}(z,v)},\quad z\in\Omega,\,v\in T_{z}(\Omega)$ (2.1)
@@
Note that (2.1) holds due to the existence of a complex geodesic $\varphi:\mathbb{D}\rightarrow\Omega$ with $\varphi(\mathbb{D})\subset V$, $\varphi(0)=z$, and $\varphi^{\prime}(0)\lambda=v$ for some $\lambda>0$, and a holomorphic map $\rho:\Omega\rightarrow\mathbb{D}$ with $\rho\circ\varphi=\text{id}_{\mathbb{D}}$. When $\Omega$ has a $C^{3}$ boundary, Theorem 2.2 was obtained by Bracci, Fornæss, and Wold [3]; in this case, the existence of $\varphi$ was given by Huang in [23, Theorem 1], and the existence of $\rho$ was given in [3, Theorem 1.1]. See also [5, 26] for related results. Gaussier and Zimmer [20] provided a proof of Theorem 2.2 for the case when $\Omega$ has a $C^{2}$ boundary. They also proved the following result (see Corollary 1.2 and its proof in [20]), which will play a crucial role in our arguments.
@@
Suppose that $M$ is a Stein manifold with $\dim_{\mathbb{C}}M\geq 2$ and $g$ is a complete Kähler metric on $M$. If there exists a compact set $\Sigma\subset M$ such that $g$ has constant negative holomorphic sectional curvature on $M\setminus\Sigma$, then $g$ has constant negative holomorphic sectional curvature on $M$ and the universal cover of $M$ is biholomorphic to a ball in $\mathbb{C}^{\dim_{\mathbb{C}}M}$.
Seshadri and Verma established a version of Theorem 2.3 for simply connected non-positively curved Kähler manifolds in [39].
@@
By the work of Graham [21, Proposition 5], the Carathéodory metric $C_{\Omega}$ is complete on $\Omega$. If $C_{\Omega}$ is a Kähler metric on $\Omega$, then by Theorem 2.3, the Carathéodory metric $C_{\Omega}$ has constant negative holomorphic sectional curvature on $\Omega$, and the universal cover of $\Omega$ is biholomorphic to the unit ball. By Theorem 1.4, the Kobayashi metric $K_{\Omega}$ is also a complete Kähler metric of constant negative holomorphic sectional curvature on $\Omega$. Since holomorphic covering maps of bounded domains are local isometries for the Kobayashi metric (which is Kähler-Einstein metric on the Euclidean ball), and both $K_{\Omega}$ and $C_{\Omega}$ are complete Kähler metrics of the same constant negative holomorphic sectional curvature on $\Omega$, we see from the Ahlfors-Schwarz lemma (see [2, 37, 44]) that $K_{\Omega}=C_{\Omega}$ on $\Omega$. By Theorem 1.3, one further concludes that $\Omega$ is biholomorphic to a ball.
@@
3 Proofs of Theorems 1.7 and 1.9
@@
where $\{\varphi_{j}\}_{j=1}^{\infty}$ is a complete orthonormal basis for the space of $L^{2}$ holomorphic functions. The Bergman metric $B_{\Omega}(z,v)$ is defined by
@@
The Bergman metric is always a Kähler metric and it is incomplete in many cases. One can also define the Bergman kernel and metric on general complex manifolds by considering $L^{2}$ holomorphic top-forms instead of functions.
@@
The remainder of the paper is dedicated to proving Theorem 1.7, Part $\bf B$, Corollary 1.8, and Theorem 1.9, which involve the Lu constant $L(\Omega)$.
@@
Since $\Omega$ is universally covered by covered by a ball, the Kobayashi metric on $\Omega$ is a Kähler metric with constant holomorphic sectional curvature. Due to the Ahlfors-Schwarz lemma (see [2, 37, 44]) and the lower bound of the holomorphic sectional curvature of the complete Bergman metric, we get

### v2

# Local Rigidity of the Bergman Metric and of the Kähler Carathéodory Metric

@@ Abstract
We prove that if the Carathéodory metric on a strictly pseudoconvex domain with a smooth boundary is locally Kähler near the boundary, then the domain is biholomorphic to a ball. We also establish a local rigidity theorem for domains with Bergman metrics of constant holomorphic sectional curvature, and highlight this relationship with the Lu constant.
footnotetext: Keywords. Carathéodory metric, Bergman metric, Kähler metric, Kähler-Einstein metric, biholomorphic map, holomorphic sectional curvature, Lu constant, strictly pseudoconvex domain, rigidity theorem Mathematics Subject Classification. Primary 32F45; Secondary 32T15, 32H02
@@ 1 Introduction
Question. What are the domains where the Bergman, Carathéodory, Kähler-Einstein, and Kobayashi metrics are not pairwise distinct (up to scaling)?
@@ 2. $2$
    where $t$ is measured with respect to the Poincaré metric on $\mathbb{D}$.
@@
Gaussier and Zimmer recently established the following beautiful relationship between local curvature conditions and global structure for complete Kähler metrics (see Corollary 1.2 and its proof in [23], and a previous result of Seshadri and Verma [44] for simply connected non-positively curved Kähler manifolds), which shows that constant negative holomorphic sectional curvature on the complement of a compact set ensures the universal cover is biholomorphic to a ball.
Theorem 1.3 ([23]).
Suppose that $M$ is a Stein manifold with $\dim_{\mathbb{C}}M\geq 2$ and $g$ is a complete Kähler metric on $M$. If there exists a compact set $\Sigma\subset M$ such that $M\backslash\Sigma$ is connected and $g$ has constant negative holomorphic sectional curvature on $M\backslash\Sigma$, then the universal cover of $M$ is biholomorphic to a ball in $\mathbb{C}^{\dim_{\mathbb{C}}M}$.
In the first main result of this paper, we discover that for Bergman metrics specifically, an even stronger phenomenon under more localized conditions - the domain itself must be biholomorphic to a ball when constant curvature holds just on an open set:
Theorem 1.4.
Let $\Omega$ be a bounded domain in $\mathbb{C}^{n}$. Suppose the Bergman metric $B^{2}_{\Omega}$ is complete and its holomorphic sectional curvature is identically equal to a negative constant $-c^{2}$ on an open set $U\subset\Omega$. Then, $\Omega$ is biholomorphic to $\mathbb{B}^{n}$, where $n=2c^{-2}-1$.
This observation advances the classical work in this area. A famous theorem of Lu [36] says that a bounded domain in $\mathbb{C}^{n}$ with a complete Bergman metric of constant holomorphic sectional curvature must be biholomorphic to $\mathbb{B}^{n}$. While the global constant curvature follows from the identity theorem of real-analytic functions, our proof provides an explicit construction of the Bergman representative coordinate $T$ as a holomorphic isometry, which has independent interest and applications. Similar uniqueness principles based on real-analyticity appear in our analysis of strictly pseudoconvex domains. Our result shows that Lu’s conclusion holds under weaker hypotheses, requiring the constant curvature condition only on an open set. In [18, 19], the first and third authors extended Lu’s theorem to the Bergman-incomplete situation and provided a multidimensional generalization of Carathéodory’s theorem on the continuous extension of biholomorphisms up to the closures. See also the recent works [28] by Huang and Li, and [21] by Ebenfelt, Treuer and Xiao.
These geometric insights naturally lead us to examine the quantitative relationships between different metrics on bounded domains. A foundational result in this direction is Lu’s inequality, which establishes precise control of the Carathéodory metric by the Bergman metric:
Theorem 1.5 ([35], cf. [10]).
$B_{\Omega}(z,v)\geq C_{\Omega}(z,v),$
The above inequality due to Lu, while fundamental, is not sharp. To obtain optimal control of the relationship between these metrics, we consider the following Lu constant.
[…]
The definition of $L(\Omega)$ implies that
$L(\Omega)B_{\Omega}\geq C_{\Omega}.$ (1.1)
Theorem 1.6.
${B}\geq 2c^{2}_{B};$
$B(z_{0})=2c^{2}_{B}(z_{0}).$
Since a bounded pseudoconvex domain with a spherical boundary in $\mathbb{C}^{n}$, $n\geq 2$, is universally covered by a ball by the uniformization theorem (see [41, Theorem A.2]), we obtain the following corollary relating to the examples of Burns and Shnider [8] using the aforementioned methods:
Corollary 1.7.
In the second half of the paper, we turn our attention to bounded domains where the Carathéodory metric exhibits special geometric properties - specifically, when it coincides up to scaling with the Bergman or Kähler-Einstein metric, or more generally, when it is a Kähler metric. This investigation connects to a broader question raised (in a slightly different form) by Yau [50, pp. 679] concerning the characterization of manifolds whose Bergman metrics are Kähler-Einstein.
On strictly pseudoconvex domains, significant progress has been made in understanding the relationships between these metrics. Cheng and Yau [11] showed the existence of complete Kähler-Einstein metrics on smoothly bounded strictly pseudoconvex domains in $\mathbb{C}^{n}$, while Mok and Yau [38] later extended this to general bounded pseudoconvex domains without boundary regularity conditions. For smoothly bounded strictly pseudoconvex domains, Cheng conjectured that the Bergman metric is Kähler-Einstein if and only if the domain is biholomorphic to a ball. Following the earlier works of Fu and the third author [22] and Nemirovski and Shafikov [41], this conjecture was confirmed by Huang and Xiao in [29]. Additionally, see the recent work [43] by Savale and Xiao for finite type domains in $\mathbb{C}^{2}$.
Building on these foundations and inspired by recent work of Gaussier and Zimmer [23], who characterized domains whose universal coverings are biholomorphic to a ball through properties of the Kobayashi metric, we establish parallel results for the Carathéodory metric. Their characterization showed:
@@ 4. $(d)$
For strictly pseudoconvex domains, we obtain the following main result, which is a Carathéodory-analogue of both Gaussier-Zimmer’s theorem and the resolution of the Cheng conjecture:
@@ 1. $(1)$
    The Carathéodory metric on $\Omega$ is a Kähler metric on $U\cap\Omega$, where $U$ is an open set that contains the boundary of $\Omega$;
@@ 1. $(i)$
    By the remarkable works of Lempert [33, 31, 34], the Carathéodory metric and the Kobayashi metric coincide on convex domains. Theorem 1.1/1.2 implies that the Carathéodory metric on a bounded convex domain is a Hermitian metric if and only if the domain is biholomorphic to the unit ball. In particular, let $\mathcal{P}$ be a small generic perturbation of $\mathbb{B}^{n}$ such that $\mathcal{P}$ is convex in $\mathbb{C}^{n}$ but not biholomorphic to $\mathbb{B}^{n}$ when $n\geq 2$ (see the papers of Burns, Shnider and Wells [9], and Greene and Krantz[26]). Then, the Carathéodory metric on $\mathcal{P}$ is not a Hermitian metric when $n\geq 2$.
@@ 3. $(iii)$
Lastly, we extend a previous result [13] (see also [12]) by Cheung and the third author from bounded convex domains to strictly pseudoconvex domains.
@@
Our proofs of Theorems 1.9 and 1.10 depend heavily on the work of Bracci, Fornæss, and Wold [5] and that of Gaussier and Zimmer [23]. (For related observations, see earlier works by Lempert [32] and Huang [27].) One of the key ideas is that at a point sufficiently close to the boundary of a smoothly bounded strictly pseudoconvex domain, there is a nonempty, open set of directions where the Kobayashi and Carathéodory metrics coincide. We observe that the Carathéodory metric has constant negative holomorphic sectional curvature in these directions, and in fact, has constant negative holomorphic sectional curvature near the boundary, if the metric is Kähler. The proof of Theorem 1.10 also relies on our Theorem 1.4.
The organization of the paper is as follows. In Section 2, we give a proof of our main result (Theorem 1.9) characterizing when the Carathéodory metric is Kähler. In Section 3, we establish our local-to-global principle for the Bergman metric (Theorem 3.4) which yields our first main theorem (Theorem 1.4), and then provide proofs of our rigidity results (Theorems 1.6 and 1.10) relating the Lu constant to geometric properties of domains.
2 When is the Carathéodory metric a Kähler metric?
@@
Theorem 2.2.
Suppose that $\Omega\subset\mathbb{C}^{n},n\geq 2,$ is a bounded strictly pseudoconvex domain with $C^{2}$ boundary. Then there exists $\epsilon>0$ such that: if $z\in\Omega$ and $\delta_{\Omega}(z)<\epsilon$, then there exists an non-empty open set $E_{\Omega}(z)$ of non-zero tangent vectors at $z$ where the Kobayashi and Carathéodory metrics agree.
Theorem 2.2 is fundamental for our analysis. The existence of complex geodesics in strictly pseudoconvex domains was first established by Lempert [32], who showed that the relevant estimates depend only on the normal curvature of the domain. Lempert’s original analysis demonstrated that since all estimates depend only on normal curvature, the results naturally extend to the $C^{2}$ case through Montel’s theorem, even though his original proof was given for the analytic case. Moreover, Huang [27] proved that certain extremal mappings near boundary points are complex geodesics (denoted here as $\varphi$). See also [7, 3] for related results. The existence of an open set where the Kobayashi and Carathéodory metrics coincide was first shown by Kosiński [30]. Later for $C^{3}$ domains, Bracci, Fornæss, and Wold [5, Theorem 1.1] established the existence of the left inverse (denoted here as $\rho$) and provided precise conditions for when the Kobayashi and Carathéodory metrics coincide near the boundary by showing
${C_{\Omega}(z,v)}={K_{\Omega}(z,v)},\quad z\in\Omega,\,v\in T_{z}(\Omega)$
@@
The proof of Theorem 2.2 was provided by Gaussier and Zimmer [23], who also obtained the following result which will play a crucial role in our arguments.
@@
Suppose that $M$ is a Stein manifold with $\dim_{\mathbb{C}}M\geq 2$, $\Sigma\subset M$ is a compact set where $M\backslash\Sigma$ is connected, and $g_{0}$ is a Hermitian metric on $M\backslash\Sigma$ which is complete at infinity. If $g_{0}$ is locally symmetric metric, then there exists a complete locally symmetric Hermitian metric $g$ on $M$ such that $g=g_{0}$ on $M\backslash\Sigma$.
Theorem 2.3 implies Theorem 1.3 since a Kähler metric with constant holomorphic sectional curvature is locally symmetric.
@@
By the work of Graham [24, Proposition 5], the Carathéodory metric $C_{\Omega}$ is complete on $\Omega$. If $C_{\Omega}$ is a Kähler metric on $U\cap\Omega$, then by Theorem 2.3, there exists a complete locally symmetric metric ${g}$ on $\Omega$ such that such that
${g}=C_{\Omega}\quad\text{on}\quad U\cap\Omega.$ (2.2)
Since the universal cover of $(\Omega,{g})$ is globally symmetric (hence homogeneous) and ${g}$ has constant negative holomorphic sectional curvature on $U\cap\Omega$, we then see that ${g}$ has constant negative holomorphic sectional curvature on $\Omega$. Thus, the universal cover of $\Omega$ is biholomorphic to the unit ball. By Theorem 1.8, the Kobayashi metric $K_{\Omega}$ is a complete Kähler metric of constant negative holomorphic sectional curvature on $\Omega$. Let $\pi:\mathbb{B}^{n}\rightarrow\Omega$ be the holomorphic covering map, which is a local isometry for $K_{\Omega}$. Hence, we have $\pi^{*}K_{\Omega}=\pi^{*}{g}$ with holomorphic sectional curvature equaling $-2$ on $\mathbb{B}^{n}$. At any point $z\in\Omega\setminus\Sigma$, it holds by (2.2) that
${C_{\Omega}(z,v)}={{g}(z,v)}={K_{\Omega}(z,v)},\quad v\in T_{z}(\Omega).$
By Theorem 1.2, one concludes that $\Omega$ is biholomorphic to a ball.
@@
From the proof of Theorem 1.9, we see that Condition $(a)$ in Theorem 1.8 can be relaxed to:
$(a^{\prime})$ The Kobayashi metric on $\Omega$ is a Kähler metric on $U\cap\Omega$, where $U$ is an open set that contains the boundary of $\Omega$.
@@
3 Local-to-global principle for the Bergman metric
@@
where $\{\varphi_{j}\}_{j=1}^{\infty}$ is a complete orthonormal basis for the space of $L^{2}$ holomorphic functions. The Bergman metric $B^{2}_{\Omega}(z,v)$ is defined by
@@
The Bergman metric is always a Kähler metric and it is incomplete in many cases. One can also define the Bergman kernel and metric on general complex manifolds by considering $L^{2}$ holomorphic top-forms instead of functions. The Bergman representative coordinate $T(z)=(w_{1},...,w_{n})^{\tau}$ relative to $p\in\Omega$ is defined as
$w_{\alpha}(z):=\sum_{j=1}^{n}g^{\bar{j}\alpha}(p)\left(K(z,p)^{-1}\left.\frac{\partial}{\partial\overline{t_{j}}}\right|_{t=p}K(z,t)-\left.\frac{\partial}{\partial\overline{t_{j}}}\right|_{t=p}\log K(t,t)\right),$ (3.1)
where $(g^{\bar{j}\alpha})=(g_{\alpha\bar{j}})^{-1}$. It is well known that $T(z)$ is holomorphic on $\Omega$ less the zero set of $K(\cdot,p)$, cf. [25].
@@
Next, we prove Theorem 1.6, Part $\bf B$ and Corollary 1.7, which involve the Lu constant $L(\Omega)$.
@@
Since $\Omega$ is holomorphically covered by a ball, the Kobayashi metric on $\Omega$ is a Kähler metric with constant holomorphic sectional curvature. Due to the Ahlfors-Schwarz lemma (see [2, 42, 49]) and the lower bound of the holomorphic sectional curvature of the complete Bergman metric, we get
@@
The remainder of the paper is dedicated to understanding domains where the Bergman metric has constant negative holomorphic sectional curvature on an open set. We first establish a fundamental result about such domains through a local-to-global principle (Theorem 3.4), which leads to our characterization theorem (Theorem 1.4).
Theorem 3.4 (Local-to-global principle for Bergman metrics).
Let $\Omega$ be a bounded domain in $\mathbb{C}^{n}$. Suppose the holomorphic sectional curvature of the Bergman metric $B^{2}_{\Omega}$ is identically equal to a negative constant $-c^{2}$ on an open set $U\subset\Omega$. Then, for any $p\in U$, the Bergman representative coordinate $T(z)=(w_{1},...,w_{n})^{\tau}$ defined by (3.1) is a bounded holomorphic map from $\Omega$ to the ball
:=$\mathcal{B}:=\{(w_{1},...,w_{n})^{\tau}:\sum_{\alpha,\beta=1}^{n}w_{\alpha}g_{\alpha\bar{\beta}}(p)\overline{w_{\beta}}<{2}{c^{-2}}\}.$
Moreover, $T:(\Omega,B^{2}_{\Omega})\to(\mathcal{B},\omega_{\mathcal{B}})$ is a holomorphic isometry, where $\omega_{\mathcal{B}}$ is the Bergman metric of $\mathcal{B}$. Consequently, the holomorphic sectional curvature of the Bergman metric $B^{2}_{\Omega}$ is identically equal to a negative constant $-c^{2}$ on $\Omega$.
Proof.
First, we assume that the Bergman metric at $p\in U$ satisfies
$g_{\alpha\bar{\beta}}(p)=\delta_{\alpha\beta},$ (3.3)
and in this case we will show that $T(z)$ defined by (3.1) is a bounded holomorphic map from $\Omega$ to the ball $\mathbb{B}^{n}:=\{w\in\mathbb{C}^{n}:|w|^{2}<{2}{c^{-2}}\}$.
At any $p\in U$, by the uniqueness of the Taylor expansion (see Bochner [4] and Lu [36, Lemma 3]), there exists a smaller neighbourhood $U_{p}$ such that the Bergman kernel can be locally decomposed as
$K(z,z)=\left(1-\frac{c^{2}}{2}|T(z)|^{2}\right)^{\frac{-2}{c^{2}}}e^{f(T(z))+\overline{f(T(z))}},\quad z\in U_{p},$ (3.4)
where $f$ is holomorphic on $U_{p}$. The map $T$ defined by (3.1) is holomorphic on $\Omega\setminus A_{p}$, where $A_{p}:=\{z\in\Omega\,|\,K(z,p)=0\,\}$ is the zero set of the Bergman kernel $K(\cdot,p)$. Let $\Omega^{\prime}:=\{z\in\Omega\setminus A_{p}:T(z)\in\mathbb{B}^{n}\}$ be the set of points in $\Omega\setminus A_{p}$ that are mapped into the ball. In particular, $U_{p}\subset\Omega^{\prime}$. By (3.4) and the theory of power series, one may duplicate the variable with its conjugate so that the full Bergman kernel can be complex analytically continued as
$K(z,{z_{0}})=\left(1-\frac{c^{2}}{2}\sum_{\alpha=1}^{n}w_{\alpha}(z)\overline{w_{\alpha}(z_{0})}\right)^{\frac{-2}{c^{2}}}e^{f(T(z))+\overline{f(T({z_{0}}))}},\quad z,z_{0}\in U_{p}.$
Consider two Kähler potentials
$\Phi_{p}(z):=\log\frac{K(z,z)K(p,p)}{|K(z,p)|^{2}},\quad z\in\Omega\setminus A_{p}$
and
$\Phi_{z_{0}}(z):=\log\frac{K(z,z)K({z_{0}},{z_{0}})}{|K(z,{z_{0}})|^{2}},\quad z\in\Omega\setminus A_{z_{0}}$
for the Bergman metric $B^{2}_{\Omega}=\partial\overline{\partial}\Phi_{z_{0}}=\partial\overline{\partial}\Phi_{p}$, for $z\in(\Omega\setminus A_{p})\cap(\Omega\setminus A_{z_{0}})$. When $U_{p}$ is small enough, for any $z_{0}\in U_{p}$, it holds that
$\Phi_{z_{0}}(z)$ $=\log\frac{\left(1-\frac{c^{2}}{2}|T(z)|^{2}\right)^{\frac{-2}{c^{2}}}e^{f(T(z))+\overline{f(T(z))}}\left(1-\frac{c^{2}}{2}|T(z_{0})|^{2}\right)^{\frac{-2}{c^{2}}}e^{f(T(z_{0}))+\overline{f(T(z_{0}))}}}{\left|1-\frac{c^{2}}{2}\sum_{\alpha=1}^{n}w_{\alpha}(z)\overline{w_{\alpha}(z_{0})}\right|^{\frac{-4}{c^{2}}}|e^{f(T(z))+\overline{f(T({z_{0}}))}}|^{2}}$
$={\frac{-2}{c^{2}}}\log\left[\left(1-\frac{c^{2}}{2}|T(z)|^{2}\right)\left(1-\frac{c^{2}}{2}|T(z_{0})|^{2}\right)\left|1-\frac{c^{2}}{2}\sum_{\alpha=1}^{n}w_{\alpha}(z)\overline{w_{\alpha}(z_{0})}\right|^{-2}\right],\quad z\in U_{p},$
which yields that
$\Phi_{p}(z_{0})=\Phi_{z_{0}}(p)={\frac{-2}{c^{2}}}\log\left(1-\frac{c^{2}}{2}|T(z_{0})|^{2}\right).$
Note that $\Phi_{p}(z)$ is defined on $\Omega\setminus A_{p}$ and thus on $\Omega^{\prime}$, where ${\frac{-2}{c^{2}}}\log\left(1-\frac{c^{2}}{2}|T(z)|^{2}\right)$ can be defined. Since these two real-analytic functions coincide on $U_{p}$, by the identity theorem they are identical to each other on $\Omega^{\prime}$. That is,
$\Phi_{p}(z)={\frac{-2}{c^{2}}}\log\left(1-\frac{c^{2}}{2}|T(z)|^{2}\right),\quad z\in\Omega^{\prime}.$ (3.5)
1) We claim that no point in $\Omega\setminus A_{p}$ is mapped outside the ball $\mathbb{B}^{n}$ by $T$.
If not, suppose there exists some point $q\in\Omega\setminus A_{p}$ that is mapped to $\{w\in\mathbb{C}^{n}:|w|^{2}\geq{2}{c^{-2}}\}$. Choose some point $q_{0}\in\Omega^{\prime}$. Since $\Omega\setminus A_{p}$ is path-connected, one can choose a path $\gamma$ that connects $q_{0}$ and $q$. Suppose under $T$ the image of $\gamma$ intersects $\partial\mathbb{B}^{n}$ firstly at some point $T(q_{1})$.
Along the path $\gamma$ take a sequence of points $(q_{l})_{l\in\mathbb{N}}\subset\Omega^{\prime}$ such that $q_{l}\to q_{1}$. Then by (3.5),
$\Phi_{p}(q_{l})={\frac{-2}{c^{2}}}\log\left(1-\frac{c^{2}}{2}|T(q_{l})|^{2}\right).$
Here, as $q_{l}\to q_{1}$, the left hand side is finite but the right hand side blows up to infinity. This is a contradiction, so we have thus proved our claim, which says that $\Omega^{\prime}=\Omega\setminus A_{p}$. Therefore, (3.5) in fact holds on $\Omega\setminus A_{p}$.
2) Since $T$ maps $\Omega\setminus A_{p}$ to the ball $\mathbb{B}^{n}$ and satisfies
$|T(z)|^{2}<{2}{c^{-2}},$ (3.6)
by the Riemann removable singularity theorem, $T$ extends across the analytic variety $A_{p}$ to the whole domain $\Omega$ with $|T(z)|^{2}\leq{2}{c^{-2}}.$ The maximum modulus principle yields that (3.6) in fact holds on $\Omega$.
Lastly, for general $p\in U$, one performs a possible linear transformation $F$ from $\Omega$ to $\Omega_{1}$ such that the Bergman metric on $\Omega_{1}$ satisfies (3.3) at $F(p)$. Since $F$ is a biholomorphism, the Bergman metric on $\Omega_{1}$ locally has the same holomorphic sectional curvature.
Denote by $\omega_{\mathcal{B}}$ the Bergman metric on $\mathcal{B}$. From the above proof, we see that the Kähler potential for the Bergman metric has the following formula.
$\Phi_{p}(z)={\frac{-2}{c^{2}}}\log\left(1-\frac{c^{2}}{2}\sum_{\alpha,\beta=1}^{n}w_{\alpha}(z)g_{\alpha\bar{\beta}}(p)\overline{w_{\beta}(z)}\right),\quad z\in\Omega.$
Direct computations then yield that $T:\Omega\to\mathcal{B}$ is an isometry. In fact, if the Bergman metric at $p\in U$ satisfies (3.3), then the potential $\Phi_{p}$ simplifies to
$\Phi_{p}(z)={\frac{-2}{c^{2}}}\log\left(1-\frac{c^{2}}{2}|T(z)|^{2}\right),\quad z\in\Omega.$
So,
$g_{\alpha\bar{\beta}}(z)$ $=\frac{\partial^{2}\Phi_{p}(z)}{\partial z_{\alpha}\partial\overline{z_{\beta}}}$
$=\sum_{i,j=1}^{n}\left(1-\frac{c^{2}}{2}|T(z)|^{2}\right)^{-2}\left[\delta_{ij}\left(1-\frac{c^{2}}{2}|T(z)|^{2}\right)+\frac{c^{2}}{2}\overline{w_{i}(z)}w_{j}(z)\right]{\frac{\partial w_{i}(z)}{\partial{z_{\alpha}}}}\frac{\partial\overline{w_{j}(z)}}{\partial{\overline{z_{\beta}}}}$
$=\sum_{i,j=1}^{n}G_{i\bar{j}}(w){\frac{\partial w_{i}(z)}{\partial{z_{\alpha}}}}\frac{\partial\overline{w_{j}(z)}}{\partial{\overline{z_{\beta}}}},\quad z\in\Omega,$
where
$G_{i\bar{j}}(w):=\frac{\partial^{2}}{\partial w_{i}\partial\overline{w_{j}}}\left(\frac{-2}{c^{2}}\log(1-\frac{c^{2}}{2}|w|^{2})\right)$
gives the Bergman metric on $\mathbb{B}^{n}$. The general case follows in a similar manner. And the curvature part also follows.
Theorem 3.4 directly implies a stronger version of Lu’s classical theorem (Theorem 1.4), requiring constant curvature only on an open set rather than globally. This insight, combined with our study of the Lu constant, provides the key to completing the proof of Theorem 1.10.